OpenStudy (anonymous):

Part 1: Using complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences? Part 2: Using these similarities and differences, how would you transform f(x) = 3 sin(4x - π) + 4 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (kohai):

I can help you with part 1 but not part 2

OpenStudy (anonymous):

Okay

OpenStudy (kohai):

So basically, the main part about the sin and cos graph is, The sin graph will always pass through (0, 0). The cos graph will not. If the graph is 2cosx, it will pass through (0, 2)

OpenStudy (kohai):

Here is what the 2cosx and 2sinx graphs look like

OpenStudy (kohai):

How high the graphs go depend on the amplitude. 2cosx 4sinx 3972478cosx The amplitudes here are 2, 4, 3972478

OpenStudy (anonymous):

Okay what about Part 2?

OpenStudy (kohai):

I can't help you with part 2, remember? :/

OpenStudy (kohai):

Oh, but this is what they'll look like

OpenStudy (kohai):

So if you graph the sin one above, all you do is move the cos one over because it can't cross the origin

OpenStudy (kohai):

Guess I could help you with part 2 c:

OpenStudy (anonymous):

So what do I write as the answer?

OpenStudy (kohai):

Draw from my explanation to answer the question

OpenStudy (anonymous):

I don't know how?

OpenStudy (anonymous):

@Hero @saifoo.khan @jdoe0001 @whpalmer4 @thomaster @SolomonZelman @e.mccormick @AccessDenied @RadEn @robtobey

OpenStudy (pink.1d):

@gamer456148

OpenStudy (pink.1d):

Part 1: The sin graph will always pass through (0, 0). The cos graph will not. If the graph is 2cosx, it will pass through (0, 2) How high the graphs go depend on the amplitude. 2cosx 4sinx 3972478cosx The amplitudes here are 2, 4, 3972478 Part 2: sin x = cos(π/2 - x) f(x) = 4 sin(2x - π) + 3 = 4 cos[π/2 - (2x - π)] + 3 = 4 cos[-2x + 3π/2] + 3 = 4 cos[-(2x - 3π/2)] + 3 = 4 cos(2x - 3π/2) + 3